OpenStudy (nottim):

How do you find the vertex of a trinomial? (e.g. x^2-10x+25) How do you find the vertex of a trinomial? (e.g. x^2-10x+25) @Mathematics

OpenStudy (anonymous):

it's easy

OpenStudy (anonymous):

Here are the steps to finding the vertex of a quadratic function: 1. Find the first derivative of the function with respect to x, and set it equal to zero. 2. Solve the first derivative for x. Set that value of x to be the x-coordinate of the vertex. 3. Now, using the original quadratic function, find the value of y at the x-coordinate of the vertex. 4. Return the tuple (x,y) corresponding to the vertex of the quadratic function.

OpenStudy (nottim):

WHAT DERIVATIVE

OpenStudy (anonymous):

Oh you don't know calculus.... one moment I will give you something else :-D

OpenStudy (anonymous):

For a trinomial \[ax^2 + bx + c\]where a, b, and c are coefficients, the x-coordinate of the vertex is \[-\frac{b}{2a}\]and the y coordinate of the vertex is equal to whatever you get when you plug in what you got for the x-coordinate. in other words: \[a(-\frac{b}{2a})^2 + b(-\frac{b}{2a}) + c\]

hero (hero):

AI, you never assume that the person knows calculus if they're asking how to find the vertex of a trinomial. What world are you living in?

hero (hero):

I'm pretty sure that if he knew calculus, he wouldn't have even asked how to find a vertex

OpenStudy (nottim):

for the x value, I've just been using the x-intercepts; adding them together and dividing it by two. what do i do for the y-value though?

OpenStudy (anonymous):

for the y-value, you simply plug into the equation what you got for the x-value.

OpenStudy (nottim):

both of them?

OpenStudy (nottim):

EVERYONE IS GONNEEE. HELP ME GHOSTS...

OpenStudy (anonymous):

just the x-value of the vertex :-P

OpenStudy (nottim):

OK

OpenStudy (anonymous):

Coordinate of vertex is (-b/2a, f(-b/2a)) 1. calculate -b/2a - it your x coordinate of vertex. 2. Plug tat value into original function instead of x - it's your y-coordinate.

hero (hero):

Not Helped

OpenStudy (anonymous):